# Explainability of Machine Learning Models under Missing Data

### v2

@@ Abstract
Missing data is a prevalent issue that can significantly impair model performance and interpretability. This paper briefly summarizes the development of the field of missing data with respect to Explainable Artificial Intelligence and experimentally investigates the effects of various imputation methods on SHAP (SHapley Additive exPlanations), a popular technique for explaining the output of complex machine learning models. Next, we compare different imputation strategies and assess their impact on feature importance and interaction as determined by Shapley values. Moreover, we also theoretically analyze the effects of missing values on Shapley values. Importantly, our findings reveal that the choice of imputation method can introduce biases that could lead to changes in the Shapley values, thereby affecting the explainability of the model. Moreover, we also show that a lower test prediction MSE (Mean Square Error) does not necessarily imply a lower MSE in Shapley values and vice versa. Also, while XGBoost (eXtreme Gradient Boosting) is a method that could handle missing data directly, using XGBoost directly on missing data can seriously affect interpretability compared to imputing the data before training XGBoost. This study provides a comprehensive evaluation of imputation methods in the context of model explanations, offering practical guidance for selecting appropriate techniques based on dataset characteristics and analysis objectives. The results underscore the importance of considering imputation effects to ensure robust and reliable insights from machine learning models.
@@ 1 Introduction
Missing data is a common issue that can significantly affect model performance and interpretability. This problem can occur due to various reasons, such as data collection errors, privacy concerns, or intentional omission. Addressing missing data is crucial, and one effective approach is imputation, i.e., filling in the missing data points to create a more complete dataset for analysis, thereby improving the overall reliability of the results.
Various imputation methods exist, ranging from simple techniques like mean imputation to more sophisticated approaches such as multiple imputation by chained equations (MICE) and k-nearest neighbors (KNN). Each method has its strengths and weaknesses, influencing not only the performance of predictive models but also their interpretability. For example, in [1], the authors studied the effects of missing data on the correlation heatmap and found out that the technique that computes a correlation matrix with the highest accuracy (in terms of root mean square error, RMSE) does not always produce correlation plots that closely resemble those derived from complete data. However, the effects of missing data on various aspects of explainable AI (XAI) have not yet been fully studied.
@@
This paper aims to explore the effects of various imputation methods on the calculation of Shapley values of the downstream classification/regression task. By comparing different imputation strategies, we seek to understand how they influence the interpretability of machine learning models and the robustness of the insights derived from Shapley values. The study provides a comprehensive evaluation of imputation methods in the context of model interpretation, offering guidance for practitioners on selecting appropriate techniques for their specific datasets and objectives.
In summary, the contributions of this paper are: (i) we explore the effects of various imputation methods on Shapley values of the downstream regression/classification task; (ii) we demonstrate that the choice of imputation method can significantly impact Shapley values, influencing the interpretability of the downstream machine learning models; (iii) we highlight that imputation methods achieving the highest prediction accuracy do not necessarily produce the most reliable feature importance values, suggesting the need for method selection aligned with dataset characteristics and analysis goals.
@@ 2.1 Explainable AI
A significant development in XAI is the emergence of intrinsic interpretability approaches, which involve designing models that are interpretable by nature. Decision trees [4], rule-based models [5], and linear models [6] are classical examples of intrinsically interpretable models. More recently, models such as Generalized Additive Models (GAMs) [7] and attention mechanisms in neural networks (NNs) [8] have been explored for their ability to provide insights into the decision-making process. However, their explanation follows their specific explanation style, making it difficult to compare with other models that do not follow the same explanation schemes during model selection.
Recent advancements have also seen the integration of XAI with deep learning models. Techniques such as Deep SHAP [9] and Integrated Gradients [10] provide explanations for neural network predictions by attributing the output to the input features in a manner consistent with the model’s internal workings. These methods bridge the gap between the high performance of deep learning models and the need for interpretability.
@@
Overall, compared to other techniques, the Shapley value offers several distinct advantages over other Explainable AI (XAI) techniques. Unlike methods that provide only local explanations or are limited to specific model architectures, the Shapley value delivers a unified and theoretically grounded approach applicable to any machine learning model. It ensures a mathematical notion of fairness by distributing the model’s output among the features based on their marginal contributions over different patterns of availability of the other features, offering an additive feature attribution that respects interaction effects. This fairness and consistency make Shapley values particularly valuable in high-stakes domains like medicine or finance [12, 13, 14, 15, 16, 17], where understanding the relative importance and interplay of various features is crucial for trust and accountability. Additionally, Shapley values provide a global interpretability framework, making it easier to understand the model’s behavior across different scenarios, thus enhancing transparency and facilitating more informed decision-making. However, the treatment of missing data can significantly influence the robustness and reliability of XAI techniques, including Shapley values.
@@ 2.2 Missing data imputation techniques
Although various missing data methods have been developed, most of them do not have a built-in explanation mechanism for imputed values. However, the way a value is imputed can have a profound impact on the performance and interpretability of downstream machine learning models. If one uses a prediction model such as a regression or classification model, then one can use Shapley values or some other XAI technique to predictions obtained from models trained after imputing training data. However, many times, one may want to use a more complicated imputation method that requires building regression or classification models and looping through the data multiple times to improve imputation accuracy. This can make it more challenging to explain the imputed values than imputing the data without using loops such as based on a regression model, for example.
@@
Attention to the explainability of imputation techniques under missing data has intensified in recent years. For example, [35] introduces Distribution-based Imputation of Missing Values with Regularization (DIMV), an interpretable imputation method that characterizes the uncertainty of imputation based on the conditional distribution of a multivariate Gaussian model, with regularization and feature selection; or [1] analyzes the effects of various imputation and direct parameter-estimation methods to the correlation plot. Next, a technique for obtaining local explanations when the point to be explained and/or the reference dataset are affected by data missingness is proposed in [36]. More specifically, they propose an extension of LIME that can be used to explain the output of models that handle missing data without imputation, such as the ones introduced in the ensuing section.
@@ 2.3 Direct missing data handling techniques without imputation
Different from the imputation approaches, methods that directly handle missing data can have clearer implications in terms of explainability. Specifically, Nguyen et al. [38] introduced an efficient Parameter Estimation for Multiple Class Monotone Missing Data (EPEM) algorithm to estimate the maximum likelihood estimates (MLEs) for multiple class monotone missing data when the covariance matrices of all classes are assumed to be equal. Additionally, Direct Parameter Estimation for Randomly missing data (DPER) [39] addresses a more general case where missing data can occur in any feature by using pairs of features to estimate the entries in the mean and covariance matrices. The implication to model explainability of using direct parameter estimation methods, like EPEM and DPER, includes improved transparency and interpretability of the model’s behavior, as these methods provide clear estimates of means and covariances, which can be directly examined and understood. In fact, [1] analyzes the effects of various imputation and direct parameter estimation methods on the correlation plot, and finds that DPER is not only scalable but also better for this task than imputation techniques such as SOFT-IMPUTE [40], KNN-based imputation (KNNI), generative adversarial imputation network (GAIN) [41], graph imputer neural network (GINN) [42], and missForest [29].
@@ 2.4 Studies on the impact of imputation on model interpretation
In the work of Shadbahr et al., it is observed that poor imputation quality can lead to misleading feature importance assessments in ML classifiers [47]. More specifically, they design a simulated dataset where they expect the distribution of the Shapley values to be symmetric, as the clusters are separated by a fixed distance and are normally distributed within those clusters. Using the skewness, they measure the symmetry of the feature importance values for each feature for all models. According to their experiments, models fit poorly imputed data giving rise to misleading feature importance metrics, demonstrating the adverse impact of poor imputation quality on model interpretability.
As highlighted by Payrovnaziri et al., changes in the underlying data due to imputation can alter model characteristics and feature importance, necessitating a careful examination of how imputed values influence predictions [48]. Their results show that the level of missingness and the imputation method used can have a significant impact on the interpretation of the models. To assess model interpretability, they focused on the stability of feature importance rankings across multiple imputations. Their findings revealed that the choice of imputation method significantly influences the interpretability of predictive models, with MICE providing more stable feature importance rankings compared to mean imputation and k-NN.
@@
The impact of missing data imputation in the performance and interpretability of prediction rule ensembles (PREs) was studied in [50]. A simulation study is conducted where a PRE model is trained using data imputed using several imputation schemes, and the model size and performance are evaluated, observing a trade-off between performance and model complexity.
@@ 6 Discussion
The variability in results between methods and missing rates underscores the need to evaluate imputation effects when using Shapley values for model interpretation for datasets with missing values. The following discussion is structured around our result and the specific pitfalls that may arise due to an incomplete understanding of the relationship between missing data, imputation methods, and Shapley values. We highlight how different approaches can lead to vastly different interpretations, how dataset characteristics and missing rates affect results, and the importance of considering both imputation accuracy and interpretability preservation:
@@
    Pitfall 1: Assume the neutrality of the imputation method. Our study reveals that different imputation methods can significantly alter Shapley values and, consequently, the interpretability of the model. For instance, mean imputation and GAIN tend to distort feature importance, especially at higher missing rates, while methods like MICE and DIMV often show similar patterns. This underscores the importance of carefully considering the imputation method when using Shapley values for model explanation, as the choice can lead to vastly different interpretations of feature importance.
@@
    Pitfall 4: Focusing solely on imputation accuracy. Our findings suggest a potential trade-off between imputation precision and the preservation of Shapley values. Methods that provide more accurate imputations do not necessarily better preserve the original Shapley values. This highlights a potential tension between optimizing for prediction accuracy and maintaining interpretability and suggests that practitioners should consider both aspects when selecting an imputation method.
These pitfalls underscore the complex interaction between missing data handling, imputation methods, and model interpretability using Shapley values. Our findings highlight that there is no universal solution for handling missing data while preserving the interpretability of the model. Instead, the choice of method should be context-dependent, considering factors such as data set characteristics, missing data rates, and the specific requirements of the analysis. Moreover, our results emphasize the need for a holistic approach that balances imputation accuracy with the preservation of feature-importance structures. As machine learning models continue to be applied in critical domains, understanding and addressing these pitfalls becomes important to ensuring reliable and interpretable results. As an additional suggestion for future work, it would be useful to develop explanation methods that are aware of whether some features in the training set have been imputed at training time and how often. Features that have been imputed more often are expected to be less reliable and, therefore, have relatively low importance.
@@ 7 Conclusion
In this paper, we explore the impact of various imputation methods on the calculation of Shapley values for model interpretability. Our findings indicate that the choice of the imputation strategy can significantly influence the accuracy and reliability of Shapley values, thereby affecting the insights drawn from the machine learning models.

### v3

@@ Abstract
Missing data is a prevalent issue that can significantly impair model performance and explainability. This paper briefly summarizes the development of the field of missing data with respect to Explainable Artificial Intelligence and experimentally investigates the effects of various imputation methods on SHAP (SHapley Additive exPlanations), a popular technique for explaining the output of complex machine learning models. Next, we compare different imputation strategies and assess their impact on feature importance and interaction as determined by Shapley values. Moreover, we also theoretically analyze the effects of missing values on Shapley values. Importantly, our findings reveal that the choice of imputation method can introduce biases that could lead to changes in the Shapley values, thereby affecting the explainability of the model. Moreover, we also show that a lower test prediction MSE (Mean Square Error) does not necessarily imply a lower MSE in Shapley values and vice versa. Also, while XGBoost (eXtreme Gradient Boosting) is a method that could handle missing data directly, using XGBoost directly on missing data can seriously affect explainability compared to imputing the data before training XGBoost. This study provides a comprehensive evaluation of imputation methods in the context of model explanations, offering practical guidance for selecting appropriate techniques based on dataset characteristics and analysis objectives. The results underscore the importance of considering imputation effects to ensure robust and reliable insights from machine learning models.
@@ 1 Introduction
Missing data is a common issue that can significantly affect model performance and explainability. This problem can occur due to various reasons, such as data collection errors, privacy concerns, or intentional omission. Addressing missing data is crucial, and one effective approach is imputation, i.e., filling in the missing data points to create a more complete dataset for analysis, thereby improving the overall reliability of the results.
Various imputation methods exist, ranging from simple techniques like mean imputation to more sophisticated approaches such as multiple imputation by chained equations (MICE) and k-nearest neighbors (KNN). Each method has its strengths and weaknesses, influencing not only the performance of predictive models but also their explainability. For example, in [1], the authors studied the effects of missing data on the correlation heatmap and found out that the technique that computes a correlation matrix with the highest accuracy (in terms of root mean square error, RMSE) does not always produce correlation plots that closely resemble those derived from complete data. However, the effects of missing data on various aspects of explainable AI (XAI) have not yet been fully studied.
@@
This paper aims to explore the effects of various imputation methods on the calculation of Shapley values of the downstream classification/regression task. By comparing different imputation strategies, we seek to understand how they influence the explainability of machine learning models and the robustness of the insights derived from Shapley values. The study provides a comprehensive evaluation of imputation methods in the context of model interpretation, offering guidance for practitioners on selecting appropriate techniques for their specific datasets and objectives.
In summary, the contributions of this paper are: (i) we explore the effects of various imputation methods on Shapley values of the downstream regression/classification task; (ii) we demonstrate that the choice of imputation method can significantly impact Shapley values, influencing the explainability of the downstream machine learning models; (iii) we highlight that imputation methods achieving the highest prediction accuracy do not necessarily produce the most reliable feature importance values, suggesting the need for method selection aligned with dataset characteristics and analysis goals.
@@ 2.1 Explainable AI
A significant development in XAI is the emergence of intrinsic explainability approaches, which involve designing models that are interpretable by nature. Decision trees [4], rule-based models [5], and linear models [6] are classic examples of intrinsically interpretable models. More recently, models such as Generalized Additive Models (GAMs) [7] and attention mechanisms in neural networks (NNs) [8] have been explored for their ability to provide insights into the decision-making process. However, their explanation follows their specific explanation style, making it difficult to compare with other models that do not follow the same explanation schemes during model selection.
Recent advancements have also seen the integration of XAI with deep learning models. Techniques such as Deep SHAP [9] and Integrated Gradients [10] provide explanations for neural network predictions by attributing the output to the input features in a manner consistent with the model’s internal workings. These methods bridge the gap between the high performance of deep learning models and the need for explainability.
@@
Overall, compared to other techniques, the Shapley value offers several distinct advantages over other Explainable AI (XAI) techniques. Unlike methods that provide only local explanations or are limited to specific model architectures, the Shapley value delivers a unified and theoretically grounded approach applicable to any machine learning model. It ensures a mathematical notion of fairness by distributing the model’s output among the features based on their marginal contributions over different patterns of availability of the other features, offering an additive feature attribution that respects interaction effects. This fairness and consistency make Shapley values particularly valuable in high-stakes domains like medicine or finance [12, 13, 14, 15, 16, 17], where understanding the relative importance and interplay of various features is crucial for trust and accountability. Additionally, Shapley values provide a global explainability framework, making it easier to understand the model’s behavior across different scenarios, thus enhancing transparency and facilitating more informed decision-making. However, the treatment of missing data can significantly influence the robustness and reliability of XAI techniques, including Shapley values.
@@ 2.2 Missing data imputation techniques
Although various missing data methods have been developed, most of them do not have a built-in explanation mechanism for imputed values. However, the way a value is imputed can have a profound impact on the performance and explainability of downstream machine learning models. If one uses a prediction model such as a regression or classification model, then one can use Shapley values or some other XAI technique to predictions obtained from models trained after imputing training data. However, many times, one may want to use a more complicated imputation method that requires building regression or classification models and looping through the data multiple times to improve imputation accuracy. This can make it more challenging to explain the imputed values than imputing the data without using loops such as based on a regression model, for example.
@@
Attention to the explainability of imputation techniques under missing data has intensified in recent years. For example, [35] introduces Distribution-based Imputation of Missing Values with Regularization (DIMV), an interpretable imputation method that characterizes the uncertainty of imputation based on the conditional distribution of a multivariate Gaussian model, with regularization and feature selection; or [1] analyzes the effects of various imputation and direct parameter-estimation methods to the correlation plot. Next, a technique for obtaining local explanations when the point to be explained and/or the reference dataset is affected by data missingness is proposed in [36]. More specifically, they propose an extension of LIME that can be used to explain the output of models that handle missing data without imputation, such as the ones introduced in the ensuing section.
@@ 2.3 Direct missing data handling techniques without imputation
Different from the imputation approaches, methods that directly handle missing data can have clearer implications in terms of explainability. Specifically, Nguyen et al. [38] introduced an efficient Parameter Estimation for Multiple Class Monotone Missing Data (EPEM) algorithm to estimate the maximum likelihood estimates (MLEs) for multiple class monotone missing data when the covariance matrices of all classes are assumed to be equal. Additionally, Direct Parameter Estimation for Randomly missing data (DPER) [39] addresses a more general case where missing data can occur in any feature by using pairs of features to estimate the entries in the mean and covariance matrices. The implication to model explainability of using direct parameter estimation methods, like EPEM and DPER, includes improved transparency and explainability of the model’s behavior, as these methods provide clear estimates of means and covariances, which can be directly examined and understood. In fact, [1] analyzes the effects of various imputation and direct parameter estimation methods on the correlation plot, and finds that DPER is not only scalable but also better for this task than imputation techniques such as SOFT-IMPUTE [40], KNN-based imputation (KNNI), generative adversarial imputation network (GAIN) [41], graph imputer neural network (GINN) [42], and missForest [29].
@@ 2.4 Studies on the impact of imputation on model interpretation
In the work of Shadbahr et al., it is observed that poor imputation quality can lead to misleading feature importance assessments in ML classifiers [47]. More specifically, they design a simulated dataset where they expect the distribution of the Shapley values to be symmetric, as the clusters are separated by a fixed distance and are normally distributed within those clusters. Using the skewness, they measure the symmetry of the feature importance values for each feature for all models. According to their experiments, models fit poorly imputed data giving rise to misleading feature importance metrics, demonstrating the adverse impact of poor imputation quality on model explainability.
As highlighted by Payrovnaziri et al., changes in the underlying data due to imputation can alter model characteristics and feature importance, necessitating a careful examination of how imputed values influence predictions [48]. Their results show that the level of missingness and the imputation method used can have a significant impact on the interpretation of the models. To assess model explainability, they focused on the stability of feature importance rankings across multiple imputations. Their findings revealed that the choice of imputation method significantly influences the explainability of predictive models, with MICE providing more stable feature importance rankings compared to mean imputation and k-NN.
@@
The impact of missing data imputation in the performance and explainability of prediction rule ensembles (PREs) was studied in [50]. A simulation study is conducted where a PRE model is trained using data imputed using several imputation schemes, and the model size and performance are evaluated, observing a trade-off between performance and model complexity.
@@ 6 Discussion
The variability in results between methods and missing rates underscores the need to evaluate imputation effects when using Shapley values for model interpretation for datasets with missing values. The following discussion is structured around our result and the specific pitfalls that may arise due to an incomplete understanding of the relationship between missing data, imputation methods, and Shapley values. We highlight how different approaches can lead to vastly different interpretations, how dataset characteristics and missing rates affect results, and the importance of considering both imputation accuracy and explainability preservation:
@@
    Pitfall 1: Assume the neutrality of the imputation method. Our study reveals that different imputation methods can significantly alter Shapley values and, consequently, the explainability of the model. For instance, mean imputation and GAIN tend to distort feature importance, especially at higher missing rates, while methods like MICE and DIMV often show similar patterns. This underscores the importance of carefully considering the imputation method when using Shapley values for model explanation, as the choice can lead to vastly different interpretations of feature importance.
@@
    Pitfall 4: Focusing solely on imputation accuracy. Our findings suggest a potential trade-off between imputation precision and the preservation of Shapley values. Methods that provide more accurate imputations do not necessarily better preserve the original Shapley values. This highlights a potential tension between optimizing for prediction accuracy and maintaining explainability and suggests that practitioners should consider both aspects when selecting an imputation method.
These pitfalls underscore the complex interaction between missing data handling, imputation methods, and model explainability using Shapley values. Our findings highlight that there is no universal solution for handling missing data while preserving the explainability of the model. Instead, the choice of method should be context-dependent, considering factors such as data set characteristics, missing data rates, and the specific requirements of the analysis. Moreover, our results emphasize the need for a holistic approach that balances imputation accuracy with the preservation of feature-importance structures. As machine learning models continue to be applied in critical domains, understanding and addressing these pitfalls becomes important to ensuring reliable and interpretable results. As an additional suggestion for future work, it would be useful to develop explanation methods that are aware of whether some features in the training set have been imputed at training time and how often. Features that have been imputed more often are expected to be less reliable and, therefore, have relatively low importance.
@@ 7 Conclusion
In this paper, we explore the impact of various imputation methods on the calculation of Shapley values for model explainability. Our findings indicate that the choice of the imputation strategy can significantly influence the accuracy and reliability of Shapley values, thereby affecting the insights drawn from the machine learning models.